# An Online Learning Approach to Prompt-based Selection of Generative Models and LLMs

### v4

@@ 1 Introduction
Example on LLMs Gemini-2.5-Flash Qwen-Plus
@@
Pass rate (%) $48.78\ (\pm 0.52)$ $\bm{82.26}\ (\pm 1.95)$
subfigureExample 1: Gemini-2.5-Flash attains higher pass rate on (Python) code completion on the HumanEval benchmark ($92.3$% versus $86.5$%) while underperforms for Java-to-C++ translation on the HumanEval-X benchmark ($48.8$% versus $82.3$%).
Example on T2I models Stable Diffusion v1.5 PixArt-$\bm{\alpha}$
@@ 3.1 Kernel Methods and Random Fourier Features
[…]
@@
[…]
@@
[…]
@@ 4 Prompt-Based Model Selection as a Contextual Bandit Problem
In this section, we introduce the framework of online prompt-based selection of generative models, which is given in Protocol 1. Let $[N]:=\{1,\cdots,N\}$ for any positive integer $N\in{\mathbb{N}}_{+}$. We denote by ${\mathcal{G}}:=[G]$ the set of (prompt-based) generative models. The evaluation proceeds in $T\in{\mathbb{N}}_{+}$ iterations.
@@
[…]
@@ 5 An Optimism-based Approach for Prompt-based Selection
[…]
@@ 5.1 The PAK-UCB Algorithm
The key component in PAK-UCB is the function Compute_UCB (lines 8-17), which outputs both the KRR estimator $\widehat{\mu}_{g}$ and an uncertainty quantifier $\widehat{\sigma}_{g}$. As the weight vector $w_{g}^{\star}$ can vary across the arms, PAK-UCB constructs the KRR dataset using prompt-score pairs from iterations where model $g$ is chosen, with the corresponding indices stored in the set $\Psi_{g}$ (line 6). The estimated score is then computed by $\widehat{s}_{g}=\widehat{\mu}_{g}+\eta\widehat{\sigma}_{g}$ (line 4), which is initially set to +$\infty$+ ∞ to ensure each model is picked at least once (lines 9-10). To provide theoretical justification for our method, we show that a variant of PAK-UCB, which is illustrated in Algorithm 3 in the Appendix, can attain a squared-root regret bound. The formal statement and the proof can be found in Appendix A.
@@ 6.2 Experiment Settings
Setup 2. Prompt-based selection of LLMs. In these experiments, we focus on prompt-based selection of large language models (LLMs) on various tasks, including Sudoku-solving and code generation (code completion and translation). In the first setup, we select LLMs to solve code generation tasks. The model is given a Python code completion problem (sampled from the first $164$ tasks in BigCodeBench (Zhuo et al., 2024)) or C++-to-Java code translation problem (sampled from the HumanEval-X benchmark (Zheng et al., 2023)). Sample prompts can be found in Figure 13 and 14 in the Appendix. Particularly, the Gemini-2.5-Flash-preview model attains a higher pass@1 rate on code completion ($55.91$% versus $13.84$%) while underperforms for C++-to-Java translation ($24.39$% versus $40.91$%) compared to the Qwen-Plus model. The rewards are binary, indicating whether the generated code can pass all the test cases. The input of the PAK-UCB, Naive-KRR, and RFF-UCB-RBF methods is the $768$-dimensional embedded prompt that is output by the RoBERTa model (for code completion) or the CodeBERT model [^8] https://huggingface.co/microsoft/codebert-base (for code translation). For LinUCB and KernelUCB baselines, we also concatenate the one-hot encoded vector of the model index to the embedded prompt. The results (Figures 4(a) and LABEL:fig:code-gen-full) show that the proposed RFF-UCB-RBF method queries from the better model conditioned to the task category (Figure LABEL:fig:llm-selection).
In the second setup, we utilize two LLMs, including o3-mini and Deepseek-Chat, to solve $9\times 9$ Sudoku problems, which contains $5$-$30$ blank entries. Sample prompt can be found in Figure 11 in the Appendix. The o3-mini model attains a higher success rate on harder Sudoku problems with more than $15$ blank entries ($70.82$% versus $50.42$%), while slightly underperforms on easier Sudoku problems ($86.85$% versus $93.6$%). The rewards are binary, indicating the solution is correct or wrong. The input of the PAK-UCB, Naive-KRR, and RFF-UCB-RBF methods is the $768$-dimensional embedded prompt that is output by the RoBERTa model [^9] https://huggingface.co/docs/transformers/en/model_doc/roberta. For LinUCB and KernelUCB baselines, we also concatenate the one-hot encoded vector of the model index to the RoBERTa-embedded prompt. The results are summarized in Figures LABEL:fig:llm-selection(b) and LABEL:fig:sudoku-full, which show that the proposed RFF-UCB-RBF algorithm outperforms all the baseline methods.
@@ Acknowledgements
[…]
@@ A.1 Regret Analysis of Sup-PAK-UCB
where $d_{\textup{eff}}$ is a data-dependent quantity defined in Lemma 5 and logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$.
Notations. To facilitate the analysis, we add an subscript $t$ to all the notations in Algorithm 3 to indicate they are quantities computed at the $t$-th iteration, i.e., $\widehat{\mu}_{g,t}^{m}$ and $\widehat{\sigma}_{g,t}^{m}$ (first appeared Line 5),$\widehat{s}_{g,t}^{m}$ and $\widetilde{s}_{g,t}^{m}$ (first appeared Line 6), $\widehat{{\mathcal{G}}}_{t}^{m}$, and $\Psi_{g,t}^{m}$. In addition, we define $g_{\star,t}:=\operatorname*{arg\,max}_{g\in{\mathcal{G}}}s_{g}(y_{t})$ as the optimal model for prompt $y_{t}$ and $\widehat{g}_{\star,t}^{m}:=\operatorname*{arg\,max}_{g\in\widehat{{\mathcal{G}}}^{m}_{t}}\widehat{s}_{g,t}^{m}$ is optimistic model at stage $m$ [^10] Note that Line 14 of Algorithm 3 is rewritten as “$\Psi_{g_{t},t+1}^{m}\leftarrow\Psi_{g_{t},t+1}^{m}\cup\{t\},\Psi_{g_{t},t+1}^{m^{\prime}}\leftarrow\Psi_{g_{t},t}^{m^{\prime}}$ for any $m^{\prime}\neq m$, and $\Psi_{g,t+1}^{m^{\prime}}\leftarrow\Psi_{g,t}^{m^{\prime}}$ for any $g\neq g_{t}$ and $m\in[M]$”. In addition, we set $\Psi_{g,t+1}^{m}\leftarrow\Psi_{g,t+1}^{m}$ for all $g\in{\mathcal{G}}$ and $m\in[M]$ in Line 8.
@@
where $d_{\textup{eff}}$ is defined therein and logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$. Plugging in Equation (6) results in
@@ A.2 Auxiliary Lemmas
Let $\Psi_{g}\subseteq[T]$ be an index set such that the set of scores $\{s_{t}:t\in\Psi_{g}\}$ are independent random variables. Then, under Assumption 1, with probability at least $1-\delta$, the quantity $\widehat{\mu}_{g}$ computed in function Compute_UCB$({\mathcal{D}},y,\Psi_{g})$ satisfies that
@@
are the summation of zero mean independent random variables, where we denote by $[\cdot]_{i}$ the $i$-th element of a vector. Further, each variable satisfies that
@@ B.1 The RFF-UCB Algorithm
[…]
@@
[…]
@@ B.2 Regret analysis of Sup-RFF-UCB
[…]
@@ B.3 Auxiliary Lemmas
where $\sigma_{p}^{2}:=\mathbb{E}_{p}[\|\omega\|^{2}]$ is the second moment of the Fourier transform of $k$ [^11] For the RBF kernel with parameter $\sigma^{2}$, i.e., $k_{\textup{RBF}}^{\sigma}(y,y^{\prime})=\exp(-\frac{1}{2\sigma^{2}}\|y-y^{\prime}\|_{2}^{2})$, we have $\sigma_{p_{\textup{RBF}}}^{2}=\frac{d}{\sigma^{2}}$. Thus, we can achieve an embedding with pointwise error no more than $\epsilon$ with probability at least $1-\delta$ as long as
@@ C.1 Results on Setups 1 and 2
Refer to caption Refer to caption 1. “a woman sitting with     her dog and a toy” 2. “the little dog stands     near the toilet in     a small bathroom” 3. “a dog on a leash     drinking water     from a water bottle” 4. “a dog laying on     the floor next to a door”
avg.CLIPScore $36.37\ (\pm 0.13)$ $\bm{37.24}\ (\pm 0.09)$
@@
Refer to caption Refer to caption 1. “a motorcycle is on     the road behind a car” 2. “two cars and a motorcycle     on a road being crossed     by a herd of elephants” 3. “a car that had run     over a red fire hydrant” 4. “a taxi driving down     a city street below     tall white buildings”
avg.CLIPScore $\bm{36.10}\ (\pm 0.06)$ $35.68\ (\pm 0.15)$
@@
Refer to caption Refer to caption 1. “the rusted out remains     of a small railway line” 2. “a skier stands in the     snow outside of a train 3. “train cars parked on a     train track near a pile     of construction material” 4. “several people on horses     with a train car     in the background”
avg.CLIPScore $\bm{35.29}\ (\pm 0.08)$ $34.25\ (\pm 0.12)$
@@
Refer to caption Refer to caption 1. “a man in red shirt     holding a baseball bat” 2. “a woman holding a     baseball bat with her     head resting on it” 3. “baseball player in the     batter’s box hitting     a baseball 4. “hind view of a baseball     player, an umpire, and     a catcher”
avg.CLIPScore $32.51\ (\pm 0.05)$ $\bm{34.30}\ (\pm 0.04)$
@@
Refer to caption Refer to caption 1. “an elephant is carrying     people across a     forested area” 2. “an elephant has a     sheet on its back 3. “a small gray elephant     standing on a beach     next to a lake” 4. “a large elephant in an     open field approaching     a vehicle”
avg.CLIPScore $\bm{36.67}\ (\pm 0.05)$ $35.08\ (\pm 0.06)$
@@
Refer to caption Refer to caption 1. “a fire hydrant in a clump     of flowering bushes” 2. “a fire hydrant on a gravel     ground with a fence     behind it” 3. “a fire hydrant that     is in the grass” 4. “a toy Ford truck next to     a fire hydrant”
avg.CLIPScore $35.11\ (\pm 0.14)$ $\bm{37.23}\ (\pm 0.05)$
@@ C.3 Results on Other Conditional Generation Tasks
Refer to caption “a person on a surfboard in the air above the water” Refer to caption “a man standing in a kitchen with a dog” Refer to caption “a bowl filled with ice cream and strawberries
@@
Refer to caption “a blurry photo of a skateboarder flying through the air” Refer to caption “a cat that is standing in the grass” Refer to caption “a blue and white bowl filled with water”

### v5

# PAK-UCB Contextual Bandit: An Online Learning Approach to Prompt-Aware Selection of Generative Models and LLMs

@@ 1 Introduction
Example on
LLMs
Gemini-2.5-Flash Qwen-Plus
@@
Pass rate (%) 48.78(±0.52)48.78\ ($\pm$ 0.52) $\bm{82.26}\ (\pm 1.95)$
subfigureExample 1: Gemini-2.5-Flash attains higher pass rate on (Python) code completion on the HumanEval benchmark (92.392.3% versus 86.586.5%) while underperforms for Java-to-C++ translation on the HumanEval-X benchmark (48.848.8% versus 82.382.3%).
Example on
T2I models
Stable Diffusion v1.5 PixArt-$\bm{\alpha}$
@@ 3.1 Kernel Methods and Random Fourier Features
Let $\phi:{\mathbb{R}}^{d}\to{\mathcal{H}}$ be a feature map from the ambient space ${\mathbb{R}}^{d}$ to the (possibly infinite-dimensional) space corresponding to the associated reproducing kernel Hilbert space (RKHS) $\mathcal{H}$}. The kernel function for RKHS $\mathcal{H}$} is defined as the inner product of the representations of the inputs as $k(y,y^{\prime}):=\langle\phi(y),\phi(y^{\prime})\rangle=\phi(y)^{\top}\phi(y^{\prime})$ for every $y,y^{\prime}\in{\mathbb{R}}^{d}$, where $\langle\cdot,\cdot\rangle$ denotes the standard inner product.
@@
where $\{w_{j}\}_{j=1}^{D}\overset{\textup{i.i.d}}{\sim}p$ and $k(y,y^{\prime})$ is approximated by the inner product $\varphi(y)^{\top}\varphi(y^{\prime})$. The resulting approximate KRR estimator is
@@
where $\widetilde{\Phi}:=[\varphi(y_{i})^{\top}]_{i=1}^{n}\in{\mathbb{R}}^{n\times 2D}$ and we denote by $\widetilde{\Phi}^{*}$ its conjugate transpose, can be computed in $O(nD^{2})$ time and $O(nD)$ memory, giving substantial computational savings if $D\ll n$. For the RBF kernel $k_{\textup{RBF}}^{\sigma}(x_{1},x_{2})=\exp(-\frac{1}{2\sigma^{2}}\|x_{1}-x_{2}\|_{2}^{2})$, the PDF $p_{\textup{RBF}}$ will be the multivariate Gaussian ${\mathcal{N}}(0,\frac{1}{\sigma^{2}}\cdot I_{d})$. [^2] We note that an alternative form of RFF is given by: $\varphi(y)=\sqrt{\frac{2}{D}}\cdot[\cos(w_{1}^{\top}y+b_{1}),\cdots,\cos(w_{D}^{\top}y+b_{D})]^{\top}$, where $\{w_{j}\}_{j=1}^{D}\overset{\textup{i.i.d}}{\sim}p$ and $\{b_{j}\}_{j=1}^{D}\overset{\textup{i.i.d}}{\sim}\textup{Unif}([0,2\pi])$. For the RBF kernel, it can be shown that  (1) has a uniformly lower variance (Sutherland & Schneider, 2015).
@@ 4 Prompt-Based Model Selection as a Contextual Bandit Problem
In this section, we introduce the framework of online prompt-based selection of generative models, which is given in Protocol 1. Let $N]:=\{1,\cdots,N\}$ for any positive integer $N\in{\mathbb{N}}_{+}$. We denote by ${\mathcal{G}}:=[G]$ the set of (prompt-based) generative models. The evaluation proceeds in $T\in{\mathbb{N}}_{+}$ iterations.
@@
1:total iterations $T\in{\mathbb{N}}_{+}$, set of generators ${\mathcal{G}}=[G]$, prompt distribution $\rho\in\Delta({\mathcal{Y}})$, score function $s:{\mathcal{Y}}\times{\mathcal{X}}\to[-1,1]$, algorithm ${\mathcal{A}}:({\mathcal{Y}}\times{\mathcal{G}}\times{\mathbb{R}})^{*}\times{\mathcal{Y}}\to\Delta({\mathcal{G}})$
@@ 5 An Optimism-based Approach for Prompt-based Selection
1:iteration $T$, generators ${\mathcal{G}}=[G]$, prompt distribution $\rho$, score function $s:{\mathcal{Y}}\times{\mathcal{X}}\to[-1,1]$, positive definite kernel $k$, regularization and exploration parameters $\alpha,\eta\geq 0$
@@ 5.1 The PAK-UCB Algorithm
The key component in PAK-UCB is the function Compute_UCB (lines 8-17), which outputs both the KRR estimator $\widehat{\mu}_{g}$ and an uncertainty quantifier $\widehat{\sigma}_{g}$. As the weight vector $w_{g}^{\star}$ can vary across the arms, PAK-UCB constructs the KRR dataset using prompt-score pairs from iterations where model $g$ is chosen, with the corresponding indices stored in the set $\Psi_{g}$ (line 6). The estimated score is then computed by $\widehat{s}_{g}=\widehat{\mu}_{g}+\eta\widehat{\sigma}_{g}$ (line 4), which is initially set to $\infty$ to ensure each model is picked at least once (lines 9-10). To provide theoretical justification for our method, we show that a variant of PAK-UCB, which is illustrated in Algorithm 3 in the Appendix, can attain a squared-root regret bound. The formal statement and the proof can be found in Appendix A.
@@ 6.2 Experiment Settings
Setup 2. Prompt-based selection of LLMs. In these experiments, we focus on prompt-based selection of large language models (LLMs) on various tasks, including Sudoku-solving and code generation (code completion and translation). In the first setup, we select LLMs to solve code generation tasks. The model is given a Python code completion problem (sampled from the first $164$ tasks in BigCodeBench (Zhuo et al., 2024)) or C++-to-Java code translation problem (sampled from the HumanEval-X benchmark (Zheng et al., 2023)). Sample prompts can be found in Figure 13 and 14 in the Appendix. Particularly, the Gemini-2.5-Flash-preview model attains a higher pass@1 rate on code completion (55.9155.91% versus 13.8413.84%) while underperforms for C++-to-Java translation (24.3924.39% versus 40.9140.91%) compared to the Qwen-Plus model. The rewards are binary, indicating whether the generated code can pass all the test cases. The input of the PAK-UCB, Naive-KRR, and RFF-UCB-RBF methods is the $768$-dimensional embedded prompt that is output by the RoBERTa model (for code completion) or the CodeBERT model [^8] https://huggingface.co/microsoft/codebert-base (for code translation). For LinUCB and KernelUCB baselines, we also concatenate the one-hot encoded vector of the model index to the embedded prompt. The results (Figures 4(a) and LABEL:fig:code-gen-full) show that the proposed RFF-UCB-RBF method queries from the better model conditioned to the task category (Figure LABEL:fig:llm-selection).
In the second setup, we utilize two LLMs, including o3-mini and Deepseek-Chat, to solve $9\times 9$ Sudoku problems, which contains 55-$30$ blank entries. Sample prompt can be found in Figure 11 in the Appendix. The o3-mini model attains a higher success rate on harder Sudoku problems with more than $15$ blank entries (70.8270.82% versus 50.4250.42%), while slightly underperforms on easier Sudoku problems (86.8586.85% versus 93.693.6%). The rewards are binary, indicating the solution is correct or wrong. The input of the PAK-UCB, Naive-KRR, and RFF-UCB-RBF methods is the $768$-dimensional embedded prompt that is output by the RoBERTa model [^9] https://huggingface.co/docs/transformers/en/model_doc/roberta. For LinUCB and KernelUCB baselines, we also concatenate the one-hot encoded vector of the model index to the RoBERTa-embedded prompt. The results are summarized in Figures LABEL:fig:llm-selection(b) and LABEL:fig:sudoku-full, which show that the proposed RFF-UCB-RBF algorithm outperforms all the baseline methods.
@@ Acknowledgements
1:total iterations $T\in{\mathbb{N}}_{+}$, set of generators ${\mathcal{G}}=[G]$, prompt distribution $\rho\in\Delta({\mathcal{Y}})$, score function $s:{\mathcal{Y}}\times{\mathcal{X}}\to[-1,1]$, positive definite kernel $k:{\mathcal{Y}}\times{\mathcal{Y}}\to{\mathbb{R}}$, regularization and exploration parameters $\alpha,\eta\geq 0$, function Compute_UCB in Algorithm 2
@@ A.1 Regret Analysis of Sup-PAK-UCB
[…]
@@
[…]
@@ A.2 Auxiliary Lemmas
[…]
@@
are the summation of zero mean independent random variables, where we denote by $\cdot]_{i}$ the $i$-th element of a vector. Further, each variable satisfies that
@@ B.1 The RFF-UCB Algorithm
1:iteration $T$, generators ${\mathcal{G}}=[G]$, prompt distribution $\rho$, score function $s:{\mathcal{Y}}\times{\mathcal{X}}\to[-1,1]$, positive definite kernel $k$, regularization and exploration parameters $\alpha,\eta\geq 0$, function Compute_UCB_RFF (Algorithm 5)
@@
7:   Draw $\omega_{1},\cdots,\omega_{D}\overset{\text{i.i.d.}}{\sim}p$
@@ B.2 Regret analysis of Sup-RFF-UCB
1:total iterations $T\in{\mathbb{N}}_{+}$, set of generators ${\mathcal{G}}=[G]$, prompt distribution $\rho\in\Delta({\mathcal{Y}})$, score function $s:{\mathcal{Y}}\times{\mathcal{X}}\to[-1,1]$, positive definite kernel $k:{\mathcal{Y}}\times{\mathcal{Y}}\to{\mathbb{R}}$, regularization and exploration parameters $\alpha,\eta\geq 0$, function Compute_UCB_RFF in Algorithm 5, parameters $D_{g,t}^{m},{\mathcal{B}}_{g,1,t}^{m}$, and ${\mathcal{B}}_{g,2,t}^{m}$
@@ B.3 Auxiliary Lemmas
where $\sigma_{p}^{2}:=\mathbb{E}_{p}[\|\omega\|^{2}]$ is the second moment of the Fourier transform of kk. [^11] For the RBF kernel with parameter $\sigma^{2}$, i.e., $k_{\textup{RBF}}^{\sigma}(y,y^{\prime})=\exp(-\frac{1}{2\sigma^{2}}\|y-y^{\prime}\|_{2}^{2})$, we have $\sigma_{p_{\textup{RBF}}}^{2}=\frac{d}{\sigma^{2}}$. Thus, we can achieve an embedding with pointwise error no more than $\epsilon$ with probability at least $1-\delta$ as long as
@@ C.1 Results on Setups 1 and 2
Refer to caption Refer to caption
1. “a woman sitting with
    her dog and a toy”
2. “the little dog stands
    near the toilet in
    a small bathroom”
3. “a dog on a leash
    drinking water
    from a water bottle”
4. “a dog laying on
    the floor next to a door”
avg.CLIPScore 36.37(±0.13)36.37\ ($\pm$ 0.13) $\bm{37.24}\ (\pm 0.09)$
@@
Refer to caption Refer to caption
1. “a motorcycle is on
    the road behind a car”
2. “two cars and a motorcycle
    on a road being crossed
    by a herd of elephants”
3. “a car that had run
    over a red fire hydrant”
4. “a taxi driving down
    a city street below
    tall white buildings”
avg.CLIPScore $\bm{36.10}\ (\pm 0.06)$ 35.68(±0.15)35.68\ ($\pm$ 0.15)
@@
Refer to caption Refer to caption
1. “the rusted out remains
    of a small railway line”
2. “a skier stands in the
    snow outside of a train
3. “train cars parked on a
    train track near a pile
    of construction material”
4. “several people on horses
    with a train car
    in the background”
avg.CLIPScore $\bm{35.29}\ (\pm 0.08)$ 34.25(±0.12)34.25\ ($\pm$ 0.12)
@@
Refer to caption Refer to caption
1. “a man in red shirt
    holding a baseball bat”
2. “a woman holding a
    baseball bat with her
    head resting on it”
3. “baseball player in the
    batter’s box hitting
    a baseball
4. “hind view of a baseball
    player, an umpire, and
    a catcher”
avg.CLIPScore 32.51(±0.05)32.51\ ($\pm$ 0.05) $\bm{34.30}\ (\pm 0.04)$
@@
Refer to caption Refer to caption
1. “an elephant is carrying
    people across a
    forested area”
2. “an elephant has a
    sheet on its back
3. “a small gray elephant
    standing on a beach
    next to a lake”
4. “a large elephant in an
    open field approaching
    a vehicle”
avg.CLIPScore $\bm{36.67}\ (\pm 0.05)$ 35.08(±0.06)35.08\ ($\pm$ 0.06)
@@
Refer to caption Refer to caption
1. “a fire hydrant in a clump
    of flowering bushes”
2. “a fire hydrant on a gravel
    ground with a fence
    behind it”
3. “a fire hydrant that
    is in the grass”
4. “a toy Ford truck next to
    a fire hydrant”
avg.CLIPScore 35.11(±0.14)35.11\ ($\pm$ 0.14) $\bm{37.23}\ (\pm 0.05)$
@@ C.3 Results on Other Conditional Generation Tasks
“a person on
a surfboard
in the air above
the water”
“a man standing
in a kitchen
with a dog”
“a bowl filled
with ice cream
and strawberries
@@
“a blurry photo of
a skateboarder flying
through the air”
“a cat that is
standing in
the grass”
“a blue and
white bowl filled
with water”